### v2

# ThinkVideo: High-Quality Reasoning Video Segmentation with Chain of Thoughts

Abstract
Reasoning Video Object Segmentation is a challenging task, which generates a mask sequence from an input video and an implicit, complex text query. Existing works probe into the problem by finetuning Multimodal Large Language Models (MLLM) for segmentation-based output, while still falling short in difficult cases on videos given temporally-sensitive queries, primarily due to the failure to integrate temporal and spatial information. In this paper, we propose ThinkVideo, a novel framework which leverages the zero-shot Chain-of-Thought (CoT) capability of MLLM to address these challenges. Specifically, ThinkVideo utilizes the CoT prompts to extract object selectivities associated with particular keyframes, then bridging the reasoning image segmentation model and SAM2 video processor to output mask sequences. The ThinkVideo framework is training-free and compatible with closed-source MLLMs, which can be applied to Reasoning Video Instance Segmentation. We further extend the framework for online video streams, where the CoT is used to update the object of interest when a better target starts to emerge and becomes visible. We conduct extensive experiments on video object segmentation with explicit and implicit queries. The results show that ThinkVideo significantly outperforms previous works in both cases, qualitatively and quantitatively. Code will be released at github.com/DanielSHKao/ThinkVideo.
Figure 1: ThinkVideo is a novel framework for Reasoning Video Segmentation utilizing the zero-shot Chain-of-Thought (CoT) capability of pretrained multimodal Large Language Models, and correctly segments the fast-moving player before, during, and after his three-point shot after CoT, given a temporally sensitive query “Which player makes a three-point shot in this basketball game” which requires both spatial and temporal reasoning. Refer to our supplemental videos for showcase demos.
1 Introduction
Reasoning Video Object Segmentation (Reasoning VOS) presents a formidable challenge in the field of computer vision, requiring the generation of a mask sequence from an input video alongside an implicit and often complex text query [82, 4]. Unlike traditional vision-language tasks that directly associate visual data with textual descriptions, Reasoning VOS requires advanced cognitive capabilities given the implicit nature of the queries. This task is particularly demanding due to the dynamic nature of video data, where factors such as temporally sensitive queries, occlusions, or rapid object movements can complicate the segmentation process.
Existing approaches of reasoning segmentation focus on finetuning the multimodal Large Language Models (MLLMs) for segmentation-based tokens. While achieving good results in Reasoning VOS, they require time-consuming finetuning process and still suffer from low accuracy when temporal reasoning is required. An implicit and time sensitive query, “Which player makes a three-point shot in this basketball game?”, for instance, is more challenging than “Which player is wearing black?” (Fig. 1). A query is time or temporally sensitive if the relevant video objects with the specified action context (e.g., three-point shot), do not appear uniformly over time but rather are concentrated at certain intervals, requiring high-level temporal reasoning to correctly find them to the satisfaction of the query requirement. Even though reasoning segmentation has achieved impressive performance on images [34, 31, 41] and videos [4], current results are far from perfect in the video domain, showcasing failure in integrating temporal information with spatial and textual context in satisfying complex and time-sensitive queries.
To address these challenges, we introduce ThinkVideo, an innovative framework designed to leverage the zero-shot Chain-of-Thought (CoT) capabilities of pre-trained MLLMs (Fig. 1). Instead of directly predicting the mask tokens, ThinkVideo breaks down the reasoning process into a sequence of “thinking” steps with multiple synthetic pairs of questions and answers. This thinking procedure will generate instance-level selectivities linked to particular keyframes, then bridging the reasoning image segmentation model with the SAM2 [60] video processor to generate coherent mask sequences. This approach not only enhances the model’s ability to reason about object relationships and movements but also allows for more accurate segmentation in complex video environments.
One of the standout features of the ThinkVideo framework is its training-free nature, making it compatible with open-source and closed-source MLLMs. We leverage the state-of-the-art Gemma3 [65] and GPT-4o [29] as the “thinking” agent in our ThinkVideo framework to generate keyframe selectivity. To evaluate the effectiveness of ThinkVideo, we conduct comprehensive experiments on referring VOS datasets (e.g., MeViS [20] and Refer-DAVIS-17 [58]) and reasoning VOS datasets (e.g., ReVOS [82] and ReasonVOS [4]). We demonstrate that ThinkVideo significantly outperforms existing referring and reasoning VOS approaches, qualitatively and quantitatively, with its more general formulation based on Reasoning Video Instance Segmentation (Reasoning VIS), which can output multiple instance-level mask sequences that satisfy the user query. In addition, we propose a variant version of ThinkVideo for online video streams, where CoT is used to update the target object when a new object starts to appear in the middle of the video that aligns better with the text query.
In conclusion, our contributions include:
  • We propose ThinkVideo, a zero-shot-CoT framework for Reasoning VIS and VOS.
  • ThinkVideo significantly outperforms existing SOTA qualitatively and quantitatively.
  • We extend ThinkVideo to process online video streams.
2 Related Work
Video segmentation. Video object segmentation (VOS) tracks objects in a given video and output corresponding mask sequences. Early VOS approaches [47, 10] rely on online first-frame finetuning, thereby constrained to specific applications because of the high time cost. To speed up inference, follow-up works adopt efficient online learning algorithm [61, 7], MRF graph inference [5], temporal CNN [81, 27], capsule routing [24], tracking [30, 68, 57, 12], embedding learning [83, 66, 13] and space-time matching [28, 54, 16]. Recent VOS methods allow various control modalities, such as categories [14, 70, 78, 89, 90, 93], points [60, 59, 26], segmentation masks [15, 17, 55, 57, 60], or text query [8, 20, 75, 62, 84, 37, 76, 48, 18], making the segmentation more flexible and interactive.
Specifically, building on the advanced development of LLM, Reasoning VOS focuses on generating mask sequences from a highly implicit and complex text query. VISA [82] first selects a keyframe and follows the core idea of LISA [34] by generating a segmentation-oriented token with finetuned image-based MLLM. Video-LISA [4] and HyperSeg [71] encode the video input in a multi-grained manner and concatenate the text query, feeding both to the LLM for segmentation-based outputs. However, none of the prior works incorporate adequate consideration of CoT reasoning ability for spatial and temporal information in Reasoning VOS. Our ThinkVideo is the first zero-shot-CoT framework compatible with pre-trained MLLM and image segmentation module, making it flexible to a boarder audience.
Chain-of-Thought MLLM. Drawing inspiration from the exceptional reasoning capabilities of Large Language Models (LLMs), Multi-modal Large Language Models (MLLMs) seek to integrate visual components to enhance multimodal perception and reasoning. These models are often employed to produce captions or respond to inquiries based on visual and textual inputs, typically achieved by techniques such as prompt engineering [42, 63, 73], cross-attention mechanism [1], Q-Former [19, 36], prompt tuning [88], linear projection [32, 39], and unified model architectures [56, 80]. These MLLMs have greatly revolutionized the traditional vision-centric tasks like image segmentation [34, 79, 87, 25, 6], object detection [87, 86, 85, 11, 46], video segmentation [82, 4, 71], and object tracking [67, 53, 94].
In addition to these advancements, CoT-based LLMs have been proposed to improve the reasoning capability [72, 69]. Existing CoT approaches tend to concentrate on inference through progressive reasoning demonstrations [91, 72, 45] or by utilizing explicit prompts that clearly outline each step [33]. However, enabling CoT capabilities for MLLMs typically undergo fine-tuning on specifically designed multimodal CoT datasets [52, 92, 43] or incorporate intricate intermediate representations [51, 64], which restricts their usability and scalability due to data accessibility and time complexity. Closer to our work, recent works, such as ThinkFirst [31] and Seg-Zero [41], have successfully incorporate CoT MLLM in reasoning image segmentation, while still falling short of “thinking” in the temporal domain. In contrast, we aim to, for the first time, abridge this gap by exploring zero-shot CoT reasoning of pre-trained MLLMs and apply them in the area of video segmentation.
3 ThinkVideo: Multi-Agent Framework
ThinkVideo is a multi-agent framework that consists of three network components: an MLLM keyframe selector $\mathcal{F}_{key}$, a reasoning image segmentation model $\mathcal{F}_{seg}$, and a video processor $\mathcal{F}_{vid}$, see Fig. 2.
Figure 2: ThinkVideo for Reasoning VIS where Reasoning VOS is a special case.
Input and Output Settings. Given an input video $x_{v}=\left\{I_{i}\in\mathbb{R}^{H\times W\times 3}\right\}_{i=1}^{T}$ and a text query $q$ that can be very complex, ThinkVideo is a Reasoning Video Instance Segmentation model $\psi_{\theta}(\cdot)$ to generate a list of non-overlapping binary mask sequences $\hat{y}=\left\{\mathcal{M}_{i}\in\mathbb{R}^{T\times H\times W}\right\}_{i=1}^{k}$:
$\hat{y}=\psi_{\theta}(x_{v},q),$ (1)
where $k$ indicates the number of object instances of interest that satisfy the requirements of input query $q$. Reasoning Video Object Segmentation (VOS) is clearly reducible to Reasoning Video Instance Segmentation (VIS) by setting $k=1$.
Each output mask sequence $\mathcal{M}_{i}=\{m_{i,t}\in\mathbb{R}^{H\times W}\}_{t=1}^{T}$ corresponds to an object instance of interest with the same number of frames and size as $x_{v}$, having the property of $m_{i,t}\cap m_{j,t}=\phi$ for $i\neq j$.
Query requirement. ThinkVideo shares a similar query setting with the first Reasoning Video Object Segmentation [82], where the input query $q$ is complex or implicit (e.g., “Please find the vehicle in the video which can accommodate at least four people.”), in contrast to the Referring VOS [62] task which takes explicit and direct query (e.g. “Please segment the red car in the video.”).
3.1 Multi-Agentic Framework
Figure 3: Illustration of the CoT Process. The MLLM agent is prompted to synthesize chain-of-thought questions and answers. This CoT process enhances the in-depth temporal and spatial understanding for the keyframe candidates. Based on the CoT result, the MLLM agent outputs an instance-level object list, which contains the object-specific keyframe and description within the frame. Refer to the supplementary materials for the detailed prompt and output examples.
MLLM Keyframe selector. The keyframe selector is an MLLM agent, e.g., GPT-4o [29] or Gemma3 [65], which generates keyframe selectivites from a list of keyframe candidates, denoted as $\{I^{c}_{i}\}$, with a CoT thinking process. The ultimate goal of this MLLM agent is to find all the object instances that meet the query requirement and select their respective keyframes, where the instance can be observed in that frame with the least effort. The output contains a list of selected object instances, each of which corresponds to an object index, a respective keyframe, and a keyframe-related object description (e.g., “the player in black jersey at the right side of the frame” ).
To achieve this goal, the agent is prompted to automatically synthesize a series of questions for each keyframe candidate and answer them one-by-one. Starting from more general information like “What can be seen in the frame?” and eventually ending with detailed questions like “Are there any other objects that meet the query requirement?”, this CoT process aims to provide in-depth spatial and temporal reasoning for the agent to find the correct objects of interest.
After analyzing all the keyframe candidates, the agent will generate an output list composed of all objects of interest and the corresponding keyframes based on its progressive question-and-answer pairs. This process can be visualized in Fig. 3.
Reasoning Image Segmentation Model. The task of the reasoning image segmentation model is to predict the key mask of each selected object instance. This agent is prompted by the selected keyframes and the keyframe-related object descriptions generated from the MLLM keyframe selector.
Video Processor. The task of the video processor is to track the selected object instances over the entire video. Each key mask generated by the reasoning image segmentation model will be tracked to produce an instance-level mask sequence.
3.2 ThinkVideo: Reasoning Video Instance Segmentation
CoT can be applied to the entire video when available for reasoning instance segmentation. Given an input video $x_{v}=\left\{I_{i}\in\mathbb{R}^{H\times W\times 3}\right\}_{i=1}^{T}$ and a text query $q$, we first uniformly sample a number of keyframe candidates $\{I^{c}_{i}\in x_{v}\}_{i=1}^{T^{\prime}}$, where $T^{\prime}=\lfloor T/\xi\rfloor$ for some $\xi<T$. Then we send these candidates and the user query $q$ to the keyframe selector $\mathcal{F}_{key}$ to generate the chain-of-thought keyframe selectivity $S$, which is a list of instance-level text descriptions $\{s_{i}\}_{i=1}^{k}$ with respect to their corresponding keyframes $\{f_{i}\}_{i=1}^{k}$ (e.g., “the red car” in $I^{c}_{2}$, “the blue motorcycle” in $I^{c}_{4}$, …), where $k$ is the predicted number of target instances. This process can be formulated as:
$I^{c}_{1:T^{\prime}}=\mathrm{Sample}\left(I_{1:T},\xi\right);$ (2)
$s_{1:k},f_{1:k}=\mathcal{F}_{key}\left(q,I^{c}_{1:T^{\prime}}\right).$ (3)
Then, we use the reasoning image segmentation model $\mathcal{F}_{seg}$ to generate per-instance key masks $\{\tilde{m}_{i}\in\mathbb{R}^{H\times W}\}$ based on the the instance-level selectivity. These key masks will be sent to the video processor $\mathcal{F}_{vid}$, i.e. SAM2 [60], to predict the per-instance mask sequences. We formulate this process as:
$\tilde{m}_{i}=\mathcal{F}_{seg}\left(s_{i},f_{i}\right);\quad\hat{\mathcal{M}}_{i}=\mathcal{F}_{vid}(\tilde{m}_{i},x_{v})\quad\mathrm{for}\quad i=1,2,...,k.$ (4)
To ensure the mask sequences are not overlapping, we adopt an optional post-processing operation:
$m_{1,t}=\hat{m}_{1,t};$ (5)
$m_{i,t}=\bigcap_{j=1}^{i-1}\neg m_{j,t}\cap\hat{m}_{i,t},$ (6)
for any $t\leq T$ and $i\leq k$, where the final output $\hat{y}\in\mathbb{R}^{T\times H\times W\times k}$ is the combination of all $m_{i,t}\in\mathbb{R}^{H\times W}$.
Figure 4: ThinkVideo for Online Reasoning Video Object Segmentation.
3.3 ThinkVideo: Online Reasoning Video Object Segmentation
Our approach can be extended to handle online video streams where future frames have yet to be observed. Refer to Fig. 4. The goal of Online Reasoning Video Object Segmentation (Online Reasoning VOS) is to generate a binary mask sequence $\hat{y}=\mathcal{M}=\{m_{t}\in\mathbb{R}^{H\times W}\}_{t=1}^{T}$ from an online input video stream $\left\{I_{i}\in\mathbb{R}^{H\times W\times 3}\right\}_{i=1}^{T}$, and text query $q$, while the model receives video frames in temporal order and makes predictions only based on frames that have already been seen. Specifically, this process can be modified from Eq. 1 as:
$m_{t}=\psi_{\theta}(I_{1:t},m_{1:t-1},q).$ (7)
The final output $\hat{y}$ can be obtained by combining predictions at all time, i.e., $m_{1},m_{2},...,m_{T}$.
The ThinkVideo framework in Online ReasoningVOS version shares a similar architecture to the original ReasoningVIS, while “asking” the keyframe selector $\mathcal{F}_{key}$ for periodic updates of the key mask. Specifically, we introduce a hyper-parameter $\xi$ for the temporal frequency of keyframe selection. We denote $I^{key}_{t}$ as the selected keyframe at time $t$, and $p_{t}$ as its corresponding frame index, respectively. When given a complex text query $q$ and an online video frame $I_{t}$, where $t=n\xi+1$ for some $n\in\mathbb{Z}$, the model sends them to $\mathcal{F}_{key}$ to justify the the keyframe selectivity $S_{t}\in\{0,1\}$, which indicates whether $I_{t}$ should be chosen as a keyframe. This process can be formulated as:
$I^{key}_{0}=\texttt{Null};\quad p_{0}=0;$ (8)
$S_{t}=\mathcal{F}_{key}\left(q,I_{t}\right);$ (9)
$I^{key}_{t}=S_{t}\cdot I_{t}+\left(1-S_{t}\right)\cdot I^{key}_{\max(t-\xi,0)};$ (10)
$p_{t}=S_{t}\cdot t+(1-S_{t})\cdot p_{\max(t-\xi,0)},$ (11)
for $t=1,\xi+1,2\xi+1,...$. The selected keyframe $I^{key}_{t}$ is then used to generate a mask control and guide the video processor for $\xi$ consecutive frames. We formulate this process as the following: $m_{t}=\mathbf{0}\mathrm{~{}if~{}}I^{key}_{t^{\prime}}=\tt{Null};$ otherwise,
$\tilde{m}_{t^{\prime}}=\mathcal{F}_{seg}\left(q,I^{key}_{t^{\prime}}\right);m_{t}=\mathcal{F}_{vid}\left(\tilde{m}_{t^{\prime}},I_{t^{\prime}:t}\right),$ (12)
where $t^{\prime}=p_{\lfloor t/\xi\rfloor\cdot\xi+1}$ for any $t\leq T$. Specifically, $t^{\prime}$ implies the frame index of the most recently selected keyframe. This framework adopts a greedy strategy to periodically update the selected keyframe when an incoming frame satisfies the query requirement, then using the keyframe to track in the following frames. When none of the previous frames is selected as a keyframe, the model outputs nothing. Eventually, we obtain the mask sequence $\hat{y}$ with all $m_{t}$’s.
Figure 5: Qualitative results. By thinking with CoT, our ThinkVideo demonstrates the temporal reasoning ability previous works fail to achieve. We compare ThinkVideo with Reasoning VOS method VideoLISA [4] and CoT Reasoning Image Segmentation methods Seg-Zero [41] and ThinkFirst [31]. Ground-truth masks are shown in the last column for reference. Zoom in for visual details.
4 Experiments
In this section, we will empirically demonstrate the effectiveness of our proposed ThinkVideo. We begin by introducing the experimental settings in Sec. 4.1, followed by showing the qualitative and quantitative results and comparation of ThinkVideo with previous works in Sec. 4.2. Lastly, we conduct ablation study to analyze the keyframe candidates for GPT-4o and CoT thinking process in Sec. 4.3.
4.1 Experimental Settings
Implementation details. We implement our MLLM keyframe selector $\mathcal{F}_{key}$ with GPT-4o [29] (version “2024-05-13”) for the original ThinkVideo and LLaVA1.5-7B [40] for the online ThinkVideo. We also conduct experiments with the open-sourced Gemma3-12B [65] in ThinkVideo for users with limited access to the OpenAI service. We adopt Seg-Zero [41] as the reasoning image segmentation model $\mathcal{F}_{seg}$ and SAM2 [60] as the video processor $\mathcal{F}_{vid}$. We use the pre-trained weights of each module from their official implementation. We set $\xi=\lfloor\frac{T-1}{8}\rfloor+1$ for our experiments (i.e., the integer value that generates at most and closest to 8 keyframe candidates), and $\xi=4$ for online experiments. GPT-4o’s temperature is set to $0.5$, and the maximum number of output tokens is $2500$. To evaluate the VOS task with the VIS version of ThinkVideo, we take the union of all the predicted instance-level mask(s) as the final output.
Datasets. We compare ThinkVideo with previous works on multiple benchmarks, including referring VOS datasets (MeViS [20] and Refer-DAVIS-2017 [58]) and reasoning VOS datasets (ReVOS [82] and ReasonVOS [4]). To evaluate reasoning VOS with temporally sensitive queries, we manually sample a subset of ReasonVOS, named as T-ReasonVOS, which contains only difficult examples involving time-sensitive queries. Refer to Appendix D for more dataset descriptions.
Evaluation metrics. We follow previous works and evaluate the ThinkVideo performance with region similarity ($\mathcal{J}$), contour accuracy ($\mathcal{F}$), and their average value ($\mathcal{J\&F}$). Specifically, $\mathcal{J}$ evaluates how well the pixels of the predicted mask match the ground truth overall using intersection over union (IoU), and $\mathcal{F}$ combines the precision and recall of boundary pixels, emphasizing the accuracy at the object boundaries.
Table 1: Referring VOS results on MeViS benchmark [20].
[…]
Traditional methods without reasoning ability
URVOS [62] 27.8 25.7 29.9
LBDT [23] 29.3 27.8 30.8
MTTR [9] 30.0 28.8 31.2
ReferFormer [77] 31.0 29.8 32.2
VLT+TC [22] 35.5 33.6 37.3
LMPM [20] 37.2 34.2 40.2
LLM-based methods with reasoning ability
LISA-7B [34] 37.2 35.1 39.4
LISA-13B [34] 37.9 35.8 40.0
TrackGPT-7B [94] 40.1 37.6 42.6
TrackGPT-13B [94] 41.2 39.2 43.1
VideoLISA (OTSA) [4] 42.3 39.4 45.2
VideoLISA (Po) [4] 44.4 41.3 47.6
VISA-7B [82] 43.5 40.7 46.3
VISA-13B [82] 44.5 41.8 47.1
ThinkVideo-LLaVA1.5-7B 52.5 49.1 55.9
ThinkVideo-Gemma3-12B 52.6 48.2 57.0
ThinkVideo-GPT-4o 60.1 56.0 64.2
Table 2: Referring VOS results on Refer-DAVIS-17 benchmark [58].
[…]
Traditional methods without reasoning ability
URVOS [62] 51.6 47.2 55.9
YOFO [35] 53.3 48.8 57.8
MLSA [74] 57.9 53.8 62.0
ReferFormer [77] 61.1 58.1 64.1
SgMg [49] 63.3 60.6 66.0
OnlineRefer [75] 64.8 61.6 67.7
LLM-based methods with reasoning ability
LISA-7B [34] 64.8 62.2 67.3
LISA-13B [34] 66.0 63.2 68.8
TrackGPT-7B [94] 63.2 59.4 67.0
TrackGPT-13B [94] 66.5 62.7 70.4
VideoLISA (OTSA) [4] 67.7 63.8 71.5
VideoLISA (Po) [4] 68.8 64.9 72.7
VISA-7B [82] 69.4 66.3 72.5
VISA-13B [82] 70.4 67.0 73.8
HyperSeg [71] 71.2 - -
ThinkVideo-LLaVA1.5-7B 73.9 70.4 77.5
ThinkVideo-Gemma3-12B 74.6 70.9 78.3
ThinkVideo-GPT-4o 79.1 75.3 82.9
Table 3: Reasoning VOS results on ReasonVOS benchmark [4]. OTSA means One-Token-Seg-All; Po means post-optimization.
[…]
Traditional methods without reasoning ability
MTTR [9] 31.1 29.1 33.1
ReferFormer [77] 32.9 30.2 35.6
SOC [44] 35.9 33.3 38.5
SgMg [49] 36.2 33.7 38.7
OnlineRefer [75] 38.7 34.6 42.9
LLM-based methods with reasoning ability
LISA [34] 31.1 29.1 33.1
VideoLISA (OTSA) [4] 45.1 43.1 47.1
VideoLISA (Po) [4] 47.5 45.1 49.9
ThinkVideo-LLaVA1.5-7B 52.0 49.5 54.5
ThinkVideo-Gemma3-12B 50.7 47.5 54.0
ThinkVideo-GPT-4o 65.5 62.4 68.7
Table 4: Reasoning VOS results on ReVOS benchmark [82]. (IT) means instruction tuning with the ReVOS training set.
[…]
Traditional methods without reasoning ability
MTTR [9] 25.5 25.1 25.9
LMPM [20] 26.4 21.2 31.7
ReferFormer [77] 28.1 26.2 29.9
LLaMA-VID [38]+LMPM 26.1 20.9 31.4
LLM-based methods with reasoning ability
LISA-13B [34] 41.6 39.8 43.5
TrackGPT-7B (IT) [94] 43.6 41.8 45.5
TrackGPT-13B (IT) [94] 45.0 43.2 46.8
VISA [82] 47.5 45.3 49.7
VISA (IT) [82] 50.9 48.8 52.9
HyperSeg [71] 55.7 53.1 58.4
ThinkVideo-LLaVA1.5-7B 46.2 43.5 48.8
ThinkVideo-Gemma3-12B 47.1 43.4 50.9
ThinkVideo-GPT-4o 55.9 52.8 59.0
4.2 Main Results
Referring VOS. We report the quantitative comparisons of ThinkVideo with previous works on the tasks of Referring VOS in Tab. 2 and Tab. 2. Online methods are highlighted in gray in these tables. Results show that both of the online and offline versions of ThinkVideo significantly outperform the state of the art. ThinkVideo achieves up to $11.3\%$ improvements over the online approaches and $15.6\%$ over the offline approaches. Notably, online ThinkVideo surpasses offline baselines in most cases, while given a more challenging task setting.
Table 5: Reasoning VOS results with temporally sensitive queries on our sampled T-ReasonVOS dataset.
[…]
LISA [34] 31.0 28.6 33.3
ThinkFirst [31]+SAM2 33.6 31.1 36.1
VideoLISA (OTSA) [4] 37.4 35.1 39.6
ThinkVideo-Gemma3-12B 50.2 46.8 53.6
ThinkVideo-GPT-4o 55.5 51.2 59.9
Table 6: Ablation study for the GPT-4o input of keyframe candidates on the ReasonVOS benchmark.
[…]
$\xi=\lfloor\frac{T-1}{4}\rfloor+1$ 66.2 63.1 69.4
$\xi=\lfloor\frac{T-1}{8}\rfloor+1$ 65.5 62.4 68.7
$\xi=\lfloor\frac{T-1}{16}\rfloor+1$ 60.0 57.1 62.9
Table 7: Ablation study for online keyframe selector on Refer-DAVIS-17. * varies depending on the OpenAI server busyness. is implemented on an RTX4090.
Model $\mathcal{J\&F}$ Avg. Time (sec.)
Qwen-2.5-VL-3B [3] 72.5 0.18
LLaVA1.5-7B [40] 73.9 0.24
GPT-4o [29] 73.2 0.27*
Reasoning VOS. We demonstrate the quantitative comparison of the ReasonVOS benchmark and ReVOS benchmark in Tab. 4 and Tab. 4, respectively, where ThinkVideo achieves up to $18\%$ than state-of-the-art offline methods and up to $13.3\%$ than state-of-the-art online methods, showcasing the better reasoning capability of “thinking” first. Implementation based on Gemma3-12B performs slightly worse results than in GPT-4o, possibly due to the capability of temporal reasoning and image understanding. We further report the qualitative performance of reasoning VOS with temporal-related queries in Fig. 5. When the temporal reasoning capability is needed, ThinkVideo achieves significantly better performance than any existing works. This is also verified in the quantitative results of T-ReasonVOS dataset, as shown in Tab. 7.
Reasoning VIS and online Reasoning VOS. We demonstrate the extended features of ThinkVideo in Fig. 6. In specific, the original ThinkVideo is able to generate instance-level mask sequences. The online version of ThinkVideo can update the object of interest during the tracking process. These extensions provide ThinkVideo with more flexibility to broader audience.
4.3 Ablation Study
Analysis of hyper-parameter $\xi$. We sample a keyframe candidate for every $\xi$ frames, then concatenate the keyframe candidates into a grid image for the GPT-4o CoT process. In Tab. 7, we compare the reasoning VOS results on the ReasonVOS benchmark [4] with different $\xi$ values. Our empirical results showcase that GPT-4o achieves comparable performance when $T^{\prime}\leq 8$ (i.e., $\xi\geq\lfloor\frac{T-1}{8}\rfloor+1$), while the image reasoning capability drastically decreases with $\xi=\lfloor\frac{T-1}{16}\rfloor+1$. Therefore, to preserve the temporal information in the video without sacrificing GPT-4o reasoning accuracy, we recommend selecting a $\xi$ value such that $4\leq T^{\prime}\leq 8$.
Analysis of CoT process. We examine the CoT process of the MLLM keyframe selector. Specifically, we show in Fig. 7 that the CoT process significantly improves the reasoning capability. In Reasoning VIS, thinking without CoT may result in missing some instances that require more effort to find. In Online Reasoning VOS, thinking without CoT degrades the performance of the keyframe justification, possibly leading to wrong classification of binary selectivity. All of these experiments point to the soundness of our proposed CoT reasoning.
Choice of online keyframe selector. We evaluate the performance of the online ThinkVideo with different keyframe selectors $\mathcal{F}_{key}$. We use the Refer-DAVIS-17 [58] benchmark for this ablation study. Tab. 7 reports the performance and average time cost of the CoT process, where LLaVA1.5-7B [40], Qwen-2.5-VL-3B [3], and GPT-4o [29] achieve comparable quantitative results. Therefore, we choose to adopt LLaVA1.5-7B in the online ThinkVideo framework.
Limitations. While ThinkVideo demonstrates strong performance in reasoning video segmentation through zero-shot Chain-of-Thought prompting, a key limitation of ThinkVideo is its reliance on pre-trained MLLMs without task-specific fine-tuning. As a result, its performance is bounded by the general reasoning capabilities of these models. We believe that fine-tuning MLLMs for temporally-aware multimodal reasoning could further improve segmentation accuracy.
Figure 6: Visual results of the original Reasoning VIS and its online version for Reasoning VOS. The original ThinkVideo can be used for instance-level reasoning video segmentation, while the online extension is preferred when the object of interest changes during the tracking process. Zoom in for details and refer to the appendix for more examples.
Figure 7: Ablation study. We examine the effectiveness of the Thinking process for keyframe selection. “Think w/o CoT” means selecting keyframes without the CoT-guided prompts. “w/o Thinking” means randomly selecting a keyframe for the original method and selecting the first frame as keyframe for its online version. Zoom in for details.
5 Conclusion
In this paper, we presented ThinkVideo, a novel zero-shot, training-free framework that leverages Chain-of-Thought (CoT) capabilities of multimodal large language models (MLLMs) to enhance reasoning in video segmentation tasks. Our extensive experiments on various benchmark datasets demonstrate that ThinkVideo significantly outperforms state-of-the-art methods, particularly in challenging scenarios involving complex, implicit and time sensitive queries, occlusions, and rapid object movements. Additionally, we extended the framework for online video streams, showcasing its adaptability for a broader audience. Overall, ThinkVideo sets a new benchmark in reasoning video segmentation and opens avenues for future research in zero-shot reasoning and its applications in dynamic visual environments.
Appendix A ThinkVideo CoT Prompt Details
One of the key components of the ThinkVideo framework is the chain-of-thought keyframe selection process using GPT-4o and Gemma3 (Eq. 3). To prompt GPT-4o [29] or Gemma3-12B [65], we first resize and concatenate all the keyframe candidates $\{I^{c}_{i}\}_{i=1}^{T^{\prime}}$ into a merged image $I^{c}_{merged}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times 3}$, where the concatenation is done on the width dimension if $W\leq H$ or on the height dimension if $H<W$, and $\max(H^{\prime},W^{\prime})=P$ for some hyperparameter $P\in\mathbb{Z}$. Then we use the following text prompt: {adjustwidth}2em2em
  Chain-of-Thought Prompt:
  ‘You will act as a keyframe selection agent for a video reasoning task. During each inference, you will be given a grid image that contains multiple keyframes sampled from a long video. The keyframes are aligned from top to down or from left to right, following their temporal order. You will also be given a complex user query that implicitly or explicitly refers to one or more target objects in the video. You need to think in chain of thoughts to analyze each keyframes and find the best keyframes for each target object, where a segmentation model can find the target object in that frame with less effort. Your chain of thoughts should begin with what can be seen in each keyframe, how many objects in total fulfilling the requirements of the user query, etc. Some of the objects may be seriously obscured or blocked by other objects. Some of the objects may be camouflaged in their surroundings. Analyze each frame separately to get all the visible objects. This chain of thoughts should follow the output format:
“Chain of Thoughts:
- Frame 1: <analysis of frame 1>;
- Frame 2: <analysis of frame 2>;
…”. For the analysis of each frame, you also have to follow the chain-of-thought format:
“- *<question 1>* <answer 1>;
- *<question 2>* <answer 2>;
…”, where you have to ask question to yourself and answer it. Your answer should be as detailed as possible. You should start with broader questions, like “what can be seen in the frame?” to some detailed questions like“are there any other objects that haven’t be listed” and “how many and which objects satisfy the user query?”. There will be many questions and answers in the analysis of each frame, helping you to fully understand the frame. The actual questions vary by cases. Generate the in-depth questions and answers based on your previous analysis. Your thinking process aims to find the keyframe for each target object of interest (find all the target objects, each of which corresponds to a keyframe) related to the user query. Lastly, you have to output a list of dictionary with a format:
“Output list: [{object__\__index: 1, keyframe: k__\__1, object__\__description: <description of the object 1 in keyframe k__\__1>}, {object__\__index: 2, keyframe: k__\__2, object__\__description: <description of the object 2 in keyframe k__\__2>}, …]”
, where each element in the list is a dictionary with three items, with object index, keyframe index, object description. k is the k-th keyframe in the grid image. object__\__index is a numbering integer starting with 1. object__\__description implies the description for that object in a particular frame, helping the model to find the object in that particular frame. For example, a valid element in an output list can be like ‘Output list: [{object__\__index: 1, keyframe: 4, object__\__description: “the man at the top left corner of the image”}]’. You have to include all objects that fulfill the requirements in the user query. Include the objects even if it is only partially visible. While choosing the keyframe for any object, you should prioritize those frames where objects are not overlapped. This will help model to better recognize the object. Keep the output list in text format. Don’t use json formatting. The output list begins with the prefix “Output list: ”, followed by a square bracket with multiple curly brackets. The square bracket should be in the same line, following the format “Output list: […]”. Don’t start with a new line.
Here is a grid image with {num__\__keyframes} keyframes. The user query is “{query}”. Follow the instruction and output the index of the best keyframe.’
[…]
Appendix B ThinkVideo (Online) CoT Prompt Details
In the online version of ThinkVideo, an incoming frame $I_{t}$ will be sent to the MLLM keyframe selector (i.e., LLaVA-7B [40]) for binary selectivity when $t=n\xi+1$. Below is the CoT prompt for our online ThinkVideo: {adjustwidth}2em2em
  Online Chain-of-Thought Prompt:
  ‘Consider the query “{query}” for an object tracking task. First analyze what can be seen in the input image and then output a simple answer with Yes or No to justify whether the input image is suitable as a keyframe. A good keyframe is an image that contains the target object. Please follow the output format “The image contains …. Therefore, the justification of using this image as keyframe is <Yes./No.>”’
  where {query} represents the user query $q$. The output of the MLLM will be parsed for binary selectivity $S_{t}$.
Appendix C Hardware Details
ThinkVideo with GPT-4o is implemented with OpenAI API and one RTX-4090 GPU. Other experiments (i.e., the ThinkVideo-Gemma3-12B and the online ThinkVideo-LLaVA1.5-7B) are implemented with two RTX-4090 GPUs.
Appendix D T-ReasonVOS Dataset
ReasonVOS contains 91 videos (18 from MeViS [20], 26 from MOSE [21], 37 from BURST  [2], and 10 from VIPSeg [50]) and 458 queries, which was proposed to effectively evaluate the performance of reasoning segmentation in videos [4]. While being widely used, this benchmark falls short in giving a fair evaluation on temporal reasoning ability. In particular, some considerable number of queries are temporally insensitive and thus less challenging, where the context only refers to the appearance or function of the relevant video objects unrelated to the actions or object movements. Thus, we believe a new dataset is both timely and essential in clearly demonstrating the efficacy of ThinkVideo on temporal reasoning. Specifically, we sampled 180 video-query pairs that correspond to 45 videos (8 from MeViS, 18 from MOSE, 16 from BURST, and 3 from VIPSeg) from ReasonVOS to produce a new benchmark called T-ReasonVOS. We remove time-insensitive queries with target objects visible from the beginning to the end of the video, such as “The mode of transportation capable of transporting the largest group of people”, where users can make decisions based on independent frames with image segmentation models. Surviving data in T-ReasonVOS after removing time-insensitive counterparts focus on videos where the target objects are visible only in a short interval, or queries specifying actions and object movements, e.g., “The thing that has made a 180-degree turn.” This requires models to “think” in the temporal dimension to find the correct objects of interest. The T-ReasonVOS dataset and the selection code will be released.
Appendix E CoT Qualitative Results
In this appendix, we will demonstrate the CoT outputs of the keyframe selector.
ThinkVideo (Original).
Refer to Appendix A for the input format. Below is two qualitative result with $T^{\prime}=7$ and $T^{\prime}=4$, respectively.
Figure A1: Input image $I^{c}_{merged}$.
  ThinkVideo Example 1 (GPT-4o):
  Input query: Please segment all the visible subjects that are using some kind of tranportation tool.
Chain of Thoughts: - Frame 1: - *What can be seen in the frame?* There are two individuals visible, both are using bicycles. - *Are any subjects using transportation tools?* Yes, both subjects are using bicycles. - *How many and which objects satisfy the user query?* There are two individuals, both satisfy the query as they are using bicycles. - *Where is the first subject located?* The first subject, wearing a white shirt, is located at the left center of the frame. - *Where is the second subject located?* The second subject, wearing a backpack, is further down the road at the center of the frame.
- Frame 2: - *What can be seen in the frame?* Three individuals are visible. Two in the foreground and one in the background, all using bicycles. - *Are any subjects using transportation tools?* Yes, all three subjects are using bicycles. - *How many and which objects satisfy the user query?* Three individuals satisfy the query using bicycles. - *Where is the first subject located?* The first subject, wearing a white shirt, is at the center left side of the frame. - *Where is the second subject located?* The second subject, in a light-colored shirt, is to the left of the first subject, occupying the left side of the frame. - *Where is the third subject located?* The third subject, with a backpack, is in the center back of the frame.
- Frame 3: - *What can be seen in the frame?* Two individuals are visible; both are on bicycles. - *Are any subjects using transportation tools?* Yes, both are using bicycles. - *How many and which objects satisfy the user query?* There are two individuals using bicycles satisfying the query. - *Where is the first subject located?* The first subject, wearing a white shirt, is centrally located in the frame. - *Where is the second subject located?* The second subject, wearing a backpack, is near the center-back of the frame.
- Frame 4: - *What can be seen in the frame?* Two individuals are visible, both using bicycles. - *Are any subjects using transportation tools?* Yes, both individuals are using bicycles as transportation. - *How many and which objects satisfy the user query?* Two subjects satisfy as they ride bicycles. - *Where is the first subject located?* The first subject, in a white shirt, is centrally situated. - *Where is the second subject located?* The second subject, with a backpack, is at the edge of the frame near the center back.
- Frame 5: - *What can be seen in the frame?* Two individuals are seen on bicycles. - *Are any subjects using transportation tools?* Yes, both are on bicycles. - *How many and which objects satisfy the user query?* Two subjects, both riding bicycles, satisfy this query. - *Where is the first subject located?* The first subject, wearing a white shirt, occupies the central area of the frame. - *Where is the second subject located?* The subject with a backpack is in the center background of the image.
- Frame 6: - *What can be seen in the frame?* Two individuals on bicycles. - *Are any subjects using transportation tools?* Yes, both are using bicycles. - *How many and which objects satisfy the user query?* Two individuals satisfy the query as they use bicycles. - *Where is the first subject located?* The first subject, in a white shirt, is placed centrally. - *Where is the second subject located?* The subject with a backpack appears in the central background.
- Frame 7: - *What can be seen in the frame?* Two individuals on bicycles. - *Are any subjects using transportation tools?* Yes, both individuals are using bicycles. - *How many and which objects satisfy the user query?* Two individuals meet the criteria, both on bicycles. - *Where is the first subject located?* The first subject, in a white shirt, is in the center. - *Where is the second subject located?* The individual with a backpack is further down, towards the center.
Output list: [{object_index: 1, keyframe: 1, object_description: “the individual in a white shirt riding a bicycle at the left center of the frame”}, {object_index: 2, keyframe: 1, object_description: “the individual with a backpack on a bicycle at the center further down the road”}, {object_index: 3, keyframe: 2, object_description: “the individual in a light-colored shirt on a bicycle at the left side of the frame”}]
Figure A2: Input image $I^{c}_{merged}$.
  ThinkVideo Example 2 (GPT-4o):
  Input query: Which item in the scene serves the purpose of drying hands after they’ve been washed?
Chain of Thoughts: - Frame 1: - *What can be seen in the frame?* A person is holding a green towel in hand, standing near a sofa. - *Is there any object that serves the purpose of drying hands?* Yes, the green towel being held by the person can serve that purpose. - *Are there any other objects that haven’t been listed?* No, the towel is the primary object fulfilling the query in this frame.
- Frame 2: - *What can be seen in the frame?* The person is standing next to a sofa, holding a green towel in both hands. - *Is the towel clearly visible?* Yes, the towel is clearly visible and being held prominently. - *Does the towel satisfy the user query?* Yes, it serves the purpose of drying hands after washing.
- Frame 3: - *What can be seen in the frame?* The person is no longer visible. The towel is now laying on the back of the sofa. - *Does any object fulfill the purpose of drying hands?* Yes, the towel on the sofa still fulfills this purpose. - *Is the towel more visible in this frame than previous frames?* The towel is distinct but lay still; it might not be as instantly recognizable as when held by the person in previous frames.
- Frame 4: - *What can be seen in the frame?* The frame is similar to Frame 3; the towel is on the back of the sofa, and no people are present. - *Does it fulfill the purpose of drying hands?* Yes, the towel on the sofa does. - *Is there any difference from Frame 3 regarding the towel’s visibility?* No significant differences are noted regarding visibility.
Output list: [{object_index: 1, keyframe: 2, object_description: “the green towel held prominently by the person”}]
ThinkVideo (Online).
Refer to Appendix B for the input format. Below is the qualitative results of two sampled frames $I_{0}$ and $I_{56}$ from the same video, where $\xi$ is set to 4:
Figure A3: Input frame $I_{0}$.
  ThinkVideo Online Example 1 (LLaVA1.5-7B):
  Input query: When parents see their children in potential danger, they are often concerned and ready to help. Who is the parent in the scene?
Chain of Thoughts: The image contains a person riding a four-wheeler in a wooded area. There is no clear indication of a parent in the scene. Therefore, the justification of using this image as a keyframe is No.
Figure A4: Input frame $I_{56}$.
  ThinkVideo Online Example 2 (LLaVA1.5-7B):
  Input query: When parents see their children in potential danger, they are often concerned and ready to help. Who is the parent in the scene?
Chain of Thoughts: The image shows the man and the child in a field, with a fence in the background. The man appears to be concerned about the child’s safety while they are on the motorcycle. Therefore, the justification of using this image as a keyframe is Yes.
Appendix F Failure Case
We demonstrate a failure case of ThinkVideo in Fig. A5. Specifically, our empirical experience shows that: when multiple object instances are visible in the video, the GPT-4o fails to identify the individual instance. Instead, similar objects are packaged within an element of the output list, e.g., “the six geese near the center of the image”. This results in the further failure of the reasoning image segmentation model. Furthermore, it is difficult to separate different instances with textual descriptions, which will be sent to the image segmentation models for each instance. Eventually, only one instance is presented in the output mask sequence in this case.
Figure A5: A failure case of ThinkVideo.
Appendix G Discussion
Keyframe selector choice. We adopt GPT-4o as the keyframe selector in the original ThinkVideo, while we also conduct experiments on additional open-sourced models for better reproducibility. We notice that smaller models, like LLaVA1.5-13B [40] or Qwen2.5-VL-3B [3], failed to follow our detailed CoT prompts. Gemma3-12B produces reasonable results and performs well in the Referring VOS datasets, while falling short in difficult reasoning tasks. More comprehensive experiments of open-sourced models on larger models, such as Llama4, remain unexplored due to the limited computational resources.
Dataset quality. We demonstrate an example from the ReasonVOS benchmark [4] of ThinkVideo in Fig. A6. Specifically, we notice that ThinkVideo segments the hoodie when asked for an object to alleviate the cold, while the ground-truth mask sequence refers to a blanket. This showcases that some of the reasoning queries are too ambiguous even for humans. While the benchmark is not our major focus and contribution, we have attempted to avoid such ambiguous queries when sampling the T-ReasonVOS dataset.
Figure A6: Ambiguous case. Zoom in for details.
Future directions. In this work, we are excited to have made significant strides in addressing the challenges of Reasoning Video Object Segmentation (Reasoning VOS) and pushing the boundaries further into Reasoning Video Instance Segmentation (Reasoning VIS) and Online Reasoning VOS. The qualitative and quantitative results of ThinkVideo are indeed impressive, showcasing its potential to enhance segmentation tasks. However, we acknowledge that our evaluation methods for Online Reasoning VIS and Reasoning VOS were not as rigorous as desired, where we only presented the qualitative results instead of quantitative results, primarily due to the limitations of available benchmarks in these tasks. This presents an opportunity for future research, where we aim to develop more comprehensive evaluation frameworks and datasets to thoroughly assess and refine the capabilities of ThinkVideo, ensuring it continues to advance the field of video segmentation.
Social impacts. Potential misuse of the ThinkVideo application could include its use in unauthorized surveillance or the creation of deceptive media. For example, individuals or organizations could leverage the technology to track people without their consent, leading to privacy violations and potential harassment. We believe that licensing can be a useful tool to prevent such issues, where we will establish specific terms and conditions governing the use of the application by the time we release the code.
Appendix H More Comparison
Fig. A7 contains three examples from T-ReasonVOS. We show that our proposed ThinkVideo consistently outperforms the state-of-the-art VideoLISA [4] on temporally sensitive queries.
Figure A7: More qualitative comparison. Refer to the anonymous website in the supplementary materials for video demo.